OpenStudy (anonymous):

What is the equation of a line passing through (3, 4) and having a slope of -5? A. y = 5x + 19 B. y = -5x + 11 C. y = -5x + 19 D. y = -1/5x +19 E. y = -1/5x+11

OpenStudy (john_es):

The slope equation of a line is, \[y=mx+n\] where m is the slope and n the y xis intercept. So m=-5, and then, \[y=-5x+n\]To find n, you know that the line must pass through (3,4), so, \[4=5\cdot3+n\Rightarrow 4=-15+n \Rightarrow n=19\]So, \[y=-5x+19\]